Khazanchi is considering an investment proposal with the following cash flows:

Initial investment - depreciable assets ........ $90,000

Initial investment - working capital......... 12,500

Net cash inflows from operations (per year for 8 years) ........ 25,000

Disinvestment - depreciable assets ...... 10,000

Disinvestment - working capital 12,500

  1. Determine the payback period.
  2. Determine the accounting rate of return in initial investment.
  3. Determine the accounting rate of return on average investement.
Answer #1

1.

Payback period refers to the time in which initial investment in the project is recovered.

Initial investment in the project = Initial investment - depreciable assets + Initial investment - working capital

= 90,000 + 12,500

= $102,500

Annual cash inflow = $25,000

Payback period = Initial investment in the project/Annual cash inflow

= 102,500/25,000

= 4.1 years

2.

Annual depreciation = (Cost of the project - Residual value)/Useful life

= (90,000 - 10,000)/8

= $10,000

Annual profit after tax = Annual cash inflow + Annual depreciation

= 25,000 + 10,000

= $35,000

Accounting rate of return = Average net income after tax/Initial investment

= 35,000/102,500

= 34.15%

3.

Average investment = [1/2 x (Initial investment in the project - Scrap value)] + Scrap value + Release of working capital

= [ 1/2 x (102,500 - 10,000)] + 10,000 + 12,500

= 46,250 + 10,000 + 12,500

= $68,750

Accounting rate of return = Average net income after tax/Average investment

= 35,000/68,750

= 50.91%